OpenStudy (anonymous):

PRETTY PLEASE HELP ME !!! Initiallly the ice cream's temperature is 0 degrees celsius. Six minutes later, it has warmed up to 8 degrees celsius. The temperature of the room is a constant 20 degrees celsius. Write an exponential function.

OpenStudy (solomonzelman):

The refrigerator doesn't work.

OpenStudy (skullpatrol):

$$\Huge y=e^{x}$$

OpenStudy (anonymous):

It is a word problem for an exponential function for melting

OpenStudy (anonymous):

@skullpatrol what would the x be

OpenStudy (skullpatrol):

The variable looks like time.

OpenStudy (anonymous):

I think SP's first answer works well, after all it doesn't ask for an exponential function that describes the situation.

OpenStudy (anonymous):

yes it does ask to use the facts to write an exponential function

OpenStudy (anonymous):

OK... 20-e^(-kx) = y (x in minutes, y in degrees) then find k using 6,8

OpenStudy (anonymous):

20-e^(-kx) = y 20-e^(-k*6) = 8 ... You got it yet?

OpenStudy (anonymous):

I do not know how to solve for that